# Monte Carlo simulations of relativistic shock breakout from a stellar wind

### v1

@@ 4 THE SPECTRUM OF ESCAPING RADIATION
Figure 4: Shock-frame, $\nu f_{\nu}$ flux of escaping photons normalized by the total kinetic energy flux of baryons at the upstream boundary, $F_{b}=\Gamma_{u}(\Gamma_{u}-1)n_{u}m_{p}c^{3}\beta_{u}$. The top, middle and bottom panels display the results for shock velocities $\Gamma_{u}=2$, $6$, and $10$, respectively. Within each panel, various lines represent different escape fractions, as specified in the legend.
As found in ILN20b for sub- and mildly relativistic shocks ($\beta_{u}=0.1$ $0.5$), the overall shape of the spectra is substantially broader than that of the Wien spectrum ($f_{\nu}\propto\nu^{3}{\rm exp}(-h\nu/kT)$). Notably, below the pronounced peak at $E_{p}$, produced by the thermal Comptonization in the downstream region, the spectra exhibit a soft extension. This extension is characterized by a slope that asymptotically approaches $f_{\nu}\propto\nu^{0}$, which is attributed to free-free emission. It extends down to energies below which the spectra show a break, resulting from the low-energy cutoff of the free-free emission due to the Coulomb screening effect. Though energetically less dominant in the spectra, the prominence of the soft tail (the spectral region where $f_{\nu}\propto\nu^{0}$) becomes more pronounced as the escape fraction increases. To clearly illustrate the feature of this soft extended tail, the spectra in the $f_{\nu}$ form are provided in Appendix A (see Figure 7). This characteristic feature, aligning with the observations for both sub-relativistic and mildly relativistic shocks as detailed in ILN20b, suggests that it can be considered a distinctive signature of shock breakout, independent of shock velocity.
A deviation from the thermal-like spectra is also observed above $E_{p}$, specifically, a non-exponential cut-off feature. While this feature is modest for $\Gamma_{u}=2$, a clear deviation becomes evident for $\Gamma_{u}=6$ and $\Gamma_{u}=10$, especially at higher escape fractions. For example, a significant excess over the exponential cutoff is observed at $f_{esc}\gtrsim 10\%$ for $\Gamma_{u}=6$ and $10$. This excess is generated by the effect of the subshock, wherein high-temperature pairs at the post-subshock region up-scatter a fraction of photons, immediately transferring their energy. Consequently, the stretch of the high-energy spectra becomes more pronounced for higher $f_{esc}$, since the post-subshock temperature increases due to the enhanced strength of the subshock as discussed in Section 3. For instance, in the simulation for $\Gamma_{u}=10$ with the highest escape fraction ($f_{esc}=41\%$), the post-subshock temperature reaches $kT\sim 4~{}{\rm MeV}$. This aligns with the observed concave curvature observed at $h\nu>E_{p}$ in the spectra, which begins to exhibit a downturn around $h\nu\sim 3kT\sim 10~{}{\rm MeV}$ (as illustrated by the red line in the bottom panel of Figure 4).
@@
Regarding $\gamma-\gamma$ attenuation, we note that high-energy photons ($h\nu>E_{p}$) undergo significant absorption before reaching the upstream boundary. For instance, in the simulation with $\Gamma_{u}=10$ and $f_{esc}=41\%$, the energy flux of counterstreaming photons at $h\nu\approx 10~{}{\rm MeV}$ at the escaping boundary ($\tau_{*}=\tau_{u}=4\times 10^{-2}$) decreases by approximately a factor of 10 compared to that at the immediate post-subshock ($\tau_{*}=0$). It is worth noting that, as will be discussed in Section 5, this $\gamma-\gamma$ attenuation of high-energy photons is not only crucial for shaping the resulting spectrum but also a dominant factor in the deceleration process of finite shocks with $\Gamma_{u}=6$ and $10$, particularly when $f_{esc}$ is large.
Additionally, particle acceleration at the subshock, which is not included in our current simulation, may have a non-negligible impact on the high-energy part of the spectrum. Specifically, pairs accelerated via shock acceleration could potentially introduce a non-thermal power-law extension in the resulting spectrum. We plan to investigate this aspect in future work.
@@ 5 Comparison with The Analytical Model
Although the model is designed to describe flows with highly relativistic velocities ($\Gamma\gg 1)$, we perform comparisons across all models, including mildly relativistic shocks with $\Gamma_{u}=2$. The comparison is conducted by fitting the Lorentz factor profile, expressed as a function of the pair-unloaded Thomson optical depth $\tilde{\tau}=2\Gamma_{u}n_{u}\sigma_{T}z$, derived from the analytical model, to the corresponding profile from the simulation through the optimization of the parameters $a$ and $\eta$ [^9] To facilitate the comparison, the optical depth used in the analytical model must be translated into $\tilde{\tau}$. The method for this translation, which is essentially identical to that adopted in GNL18 with a minor modification, is provided in Appendix C. These values depend on the energy and angular distributions of pairs and photons and are expected to be on the order of unity. Accordingly, we conduct the fitting procedure within predefined ranges, with $a$ set to vary between 0.5 and 3, and $\eta$ from 0.25 to 1.0.101010Even if we succeed in obtaining a good fit with values far beyond the predefined range, the solution should be considered unphysical. To determine the best-fit values, we systematically scanned these ranges in increments of $0.05$ for both parameters. The comparison is displayed in Figure 5. It is worth noting that $\tilde{\tau}$ represents the pair-unloaded optical depth measured in the upstream rest frame. Therefore, when a shock propagates in a wind-like environment, the fraction of shock energy released at a given optical depth can be determined by identifying $f_{esc}$ that yields a shock structure with a depth $\tilde{\tau_{u}}(=2\Gamma_{u}\sigma_{T}z_{u})$ comparable to the optical depth.
@@
Figure 5: Comparison of the Lorentz factor profile as a function of pair-unloaded optical depth, $\tilde{\tau}=2\Gamma_{u}n_{u}\sigma_{T}z$, with the analytical model of GNL18 for $\Gamma_{u}=2$ (top), $6$ (middle) and $10$ (bottom). For the shock with $\Gamma_{u}=2$, the analytical model reproduces the numerical results, with the optimal parameter pairs $(a,\eta)$ for each escape fraction $f_{esc}$ are determined as follows: at $f_{esc}=0\%$, $a=0.55$ and $\eta=0.4$; at $0.55~{}\%$, $a=0.6$ and $\eta=0.4$; at $f_{esc}=1.5~{}\%$, $a=0.65$ and $\eta=0.4$; at $f_{esc}=6~{}\%$ $a=0.7$ and $\eta=0.4$; at $f_{esc}=13~{}\%$ $a=0.8$ and $\eta=0.4$; at $f_{esc}=35~{}\%$ $a=1.95$ and $\eta=0.35$. Although the best fit values of the free parameters vary among the different escape fractions, a single pair, e.g., $a=0.5$ and $\eta=0.3$ exhibits a good agreemement with the numerical results. On the other hand, for the shock with $\Gamma_{u}=6$, the optimal parameter pairs $(a,\eta)$ for each escape fraction $f_{esc}$ are determined as follows: at $f_{esc}=0\%$, $a=2.4$ and $\eta=0.6$; at $f_{esc}=0.65\%$, $a=0.7$ and $\eta=0.75$; at $f_{esc}=2.9\%$, $a=0.5$ and $\eta=0.35$; and for $6.8\%$, $13\%$, and $46\%$, both parameters consistently reach the lower bounds of the fitting range, $a=0.5$ and $\eta=0.25$, indicating the analytical model’s limitations in capturing the simulation results in these cases. For the shock with $\Gamma_{u}=10$, the optimal pairs are as follows: at $f_{esc}=0\%$, $a=1.75$ and $\eta=0.6$; at $f_{esc}=0.41\%$, $a=0.5$ and $\eta=0.8$; $f_{esc}=2.5\%$, $a=0.5$ and $\eta=0.35$; and for $5.9\%$, $14\%$, and $41\%$, both parameters again consistently reach the lower bounds, $a=0.5$ and $\eta=0.25$, highlighting similar limitations of the analytical model in these cases.
@@ 5.1 Comparison of Shock Profiles
In the case of infinite shocks, we find that, within a reasonable range of $\eta$ and $a$, the analytical model reproduces the numerical results quite well in all cases ($\Gamma_{u}=2$, $6$, and $10$). This agreement was previously reported in ILN20a for $\Gamma_{u}=6$ and $10$ (as well as for $\Gamma_{u}=20$, which is not included in the present paper), confirming the model’s validity for highly relativistic shocks. Here, we extend this finding to mildly relativistic shocks, showing that even for $\Gamma_{u}=2$, the analytical model exhibits good agreement with the numerical results, demonstrating its applicability to lower-velocity shocks than previously expected. A more comprehensive comparison between the analytical model and the numerical results for infinite shocks is presented in Appendix C. For mildly relativistic shocks ($\Gamma_{u}=2$), we also find a good match in the case of finite shocks, regardless of the value of $f_{esc}$. On the other hand, for higher Lorentz factor shocks ($\Gamma_{u}=6$ and $10$), while good agreement persists when the escape fraction is small ($f_{esc}\lesssim 1\%$), a significant discrepancy emerges as $f_{esc}$ increases. This discrepancy cannot be reconciled within a reasonable range of the parameters $\eta$ and $a$.
@@
The result is counterintuitive, as one would naively expect highly relativistic shocks to match better with the analytical model, given that it is constructed under the assumption of $\Gamma\gg 1$. This discrepancy can be understood as a consequence of the role of the subshock in highly relativistic shocks, which influences the spectral shape of counterstreaming photons. As shown in Section 3, the subshock produces a high-temperature spike in the immediate downstream region, where relativistically heated pairs scatter photons to higher energies. While the post-subshock temperature remains in the range $kT_{d,sub}\lesssim 0.5m_{e}c^{2}$ for $\Gamma_{u}=2$ (at least up to $f_{esc}\leq 35\%$), it exceeds $m_{e}c^{2}$ and continues to rise as the escape fraction increases for $\Gamma_{u}=6$ and $10$. As a result of this high temperature, a high-energy photon component ($h\nu\sim m_{e}c^{2}$) emerges in addition to the bulk of photons with energy $\sim m_{e}c^{2}$, as described in Section 4. The high-energy extension observed in the escaping photon spectrum above the spectral peak $E_{p}\sim m_{e}c^{2}$ (Figure 4) corresponds to the fraction of these high-energy photons that counterstream up to the upstream boundary. As mentioned in Section 4, these high-energy photons are efficiently absorbed via $\gamma$-$\gamma$ attenuation before reaching the upstream boundary. This absorption of high-energy counterstreaming photons is responsible for the deviation of the numerical results from the analytical model, which can be explained as follows.
@@
The key point is that the absorption of counterstreaming high-energy photons primarily occurs via collisions with the bulk population of lower-energy photons that are likewise counterstreaming upstream from the immediate downstream region, rather than with photons advecting with the flow (i.e., backscattered photons). This occurs because the high-energy photons possess sufficient energy to exceed the pair production threshold necessary for interaction with the bulk of the counterstreaming photon population. As a result, the effective target photon population for the high-energy photons is significantly more abundant than the backscattered photons assumed in the analytical model. As the escape fraction increases, the subshock strengthens, and this channel of $\gamma-\gamma$ pair production becomes the primary mechanism governing the shock deceleration process. Consequently, as $f_{\rm esc}$ increases, the assumption in the analytical model that shock dissipation is governed by the interaction of counterstreaming photons ($h\nu\sim m_{e}c^{2}$) with pairs and backscattered photons begins to break down, and the shock deceleration width becomes significantly shorter due to the high opacity arising from $\gamma$$\gamma$ interactions among counterstreaming photons. Shocks with $\Gamma_{u}=10$ exhibit a relatively larger deviation from the analytical model compared to those with $\Gamma_{u}=6$, primarily because the temperature spike at the subshock tends to be more pronounced for larger $\Gamma_{u}$.
@@ 5.2 Comparison of Shock Widths
In Figure 6, we present the deceleration length of finite shocks, $\Delta\tilde{\tau}$, as a function of $f_{esc}$, compared with predictions from the analytical model. For the free parameters of the analytical model, we adopt the pair of $a$ and $\eta$ that best fit the $\Gamma-\tilde{\tau}$ profile of the numerical results for the lowest $f_{esc}$ cases (profiles shown by green lines in Figure 5). It is worth noting that although the analytical model considered in this study is modified from the original by shifting the reference position of $\tilde{\tau}=0$ to the location where $\Gamma=\Gamma_{u,sub}$—rather than $\Gamma=1$, as in the original analytical solution—this adjustment, while shortening the shock width by removing part of the upstream region, has a negligible impact on its overall value. This is because the shock width is primarily determined by the far upstream region, where $\Gamma\sim\Gamma_{u}$, as evident from Figure 5.
@@
Figure 6: Comparison of the shock width, $\Delta\tilde{\tau}$, obtained from the simulation (blue circles) and the analytical model (black dashed lines) as a function of $f_{esc}$ for $\Gamma_{u}=2$ (top), $6$ (middle), and $\Gamma_{u}=10$ (bottom). The shock width, $\Delta\tilde{\tau}$, is defined as the optical depth measured from the subshock to the upstream position where the flow’s four-velocity reaches 90% of its far-upstream value ($\Gamma\beta=0.9\Gamma_{u}\beta_{u}$). For the analytical model, fixed values are adopted for the pair of parameters ($a$ and $\eta$) that best fit the simulation with the lowest $f_{esc}$. The orange solid lines in the middle and bottom panels represent the best-fit function of the form $\Delta\tilde{\tau}=Cf_{esc}^{X}$, fitted to the numerical results for the $f_{esc}$$\tilde{\tau}$ trend.
@@
Since the shock profile is well reproduced by the analytical model, the dependence of the shock width on $f_{esc}$ observed in the simulation is also well captured by the analytical model for $\Gamma_{u}=2$. On the other hand, reflecting the breakdown of the analytical model at high escape fractions, the $\tilde{\tau}-f_{esc}$ dependence observed in the simulation cannot be reproduced by the analytical model for $\Gamma_{u}=6$ and $10$. Note that varying the free parameters does not alleviate this discrepancy, as evidenced by the fact that large values of $f_{esc}$ cannot be described within the framework of the analytical model, as mentioned above. At large escape fractions ($f_{esc}\gg\Gamma_{u}^{-2}$), the analytical model predicts a shock width approximately given by $\Delta\tilde{\tau}\sim\frac{m_{e}}{m_{p}}\frac{\Gamma_{u}}{f_{esc}}=5.4\times 10^{-3}\Gamma_{u,1}f_{esc}^{-1}.$ In contrast, fitting the simulation data to a power-law function of the form $\Delta\tilde{\tau}\propto f_{esc}^{X}$ yields a scaling of $\Delta\tilde{\tau}\propto f_{esc}^{-1.45}$ ($f_{esc}^{-1.63}$) for $\Gamma_{u}=6$ ($10$), exhibiting a steeper dependence on $f_{esc}$. Thus, the simulation exhibits a much narrower shock width at large $f_{esc}$, with the gap between the simulation and the analytical model exceeding an order of magnitude at the highest escape fractions.

### v2

@@ 4 THE SPECTRUM OF ESCAPING RADIATION
Figure 5: Shock-frame, $\nu f_{\nu}$ flux of escaping photons normalized by the total kinetic energy flux of baryons at the upstream boundary, $F_{b}=\Gamma_{u}(\Gamma_{u}-1)n_{u}m_{p}c^{3}\beta_{u}$. Each panel shows results for $\Gamma_{u}=2$ (top left), $3.5$ (top right), 66 (bottom left), and $10$ (bottom right).
Lines indicate different escape fractions (see legends).
As found in ILN20b for sub- and mildly relativistic shocks ($\beta_{u}=0.1$ - $0.5$), the overall shape of the spectra is substantially broader than that of the Wien spectrum ($f_{\nu}\propto\nu^{3}{\rm exp}(-h\nu/kT)$). Notably, below the pronounced peak at $E_{p}$, produced by the thermal Comptonization in the downstream region, the spectra exhibit a soft extension. This extension is characterized by a slope that asymptotically approaches $f_{\nu}\propto\nu^{0}$, which is attributed to free-free emission. It extends down to energies below which the spectra show a break, resulting from the low-energy cutoff of the free-free emission due to the Coulomb screening effect. Though energetically less dominant in the spectra, the prominence of the soft tail (the spectral region where $f_{\nu}\propto\nu^{0}$) becomes more pronounced as the escape fraction increases. To clearly illustrate the feature of this soft extended tail, the spectra in the $f_{\nu}$ form are provided in Appendix A (see Figure 8). This characteristic feature, aligning with the observations for both sub-relativistic and mildly relativistic shocks as detailed in ILN20b, suggests that it can be considered a distinctive signature of shock breakout, independent of shock velocity.
A deviation from the thermal-like spectra is also observed above $E_{p}$, specifically, a non-exponential cut-off feature. While this feature is modest for $\Gamma_{u}=2$, a clear deviation becomes evident for $\Gamma_{u}\geq 3.5$, especially at higher escape fractions. For example, a significant excess over the exponential cutoff is observed at $f_{esc}=49\%$ for $\Gamma_{u}=3.5$, and at $f_{esc}\gtrsim 10\%$ $\Gamma_{u}=6$ and $10$. This excess is generated by the effect of the subshock, wherein high-temperature pairs at the post-subshock region up-scatter a fraction of photons, immediately transferring their energy. Consequently, the stretch of the high-energy spectra becomes more pronounced for higher $f_{esc}$, since the post-subshock temperature increases due to the enhanced strength of the subshock as discussed in Section 3. For instance, in the simulation for $\Gamma_{u}=10$ with the highest escape fraction ($f_{esc}=41\%$), the post-subshock temperature reaches $kT\sim 4~{\rm MeV}$. This aligns with the observed concave curvature observed at $h\nu>E_{p}$ in the spectra, which begins to exhibit a downturn around $h\nu\sim 3kT\sim 10~{\rm MeV}$ (as illustrated by the red line in the bottom panel of Figure 5).
@@
Regarding $\gamma-\gamma$ attenuation, we note that high-energy photons ($h\nu>E_{p}$) undergo significant absorption before reaching the upstream boundary. For instance, in the simulation with $\Gamma_{u}=10$ and $f_{esc}=41\%$, the energy flux of counterstreaming photons at $h\nu\approx 10~{\rm MeV}$ at the escaping boundary ($\tau_{*}=\tau_{u}=4\times 10^{-2}$) decreases by approximately a factor of 10 compared to that at the immediate post-subshock ($\tau_{*}=0$). It is worth noting that, as will be discussed in Section 5, this $\gamma-\gamma$ attenuation of high-energy photons is not only crucial for shaping the resulting spectrum but also a dominant factor in the deceleration process of finite shocks with $\Gamma_{u}=3.5$, $6$ and $10$, particularly when $f_{esc}$ is large. Additionally, particle acceleration at the subshock—–omitted in our present simulations—–may influence the high-energy end of the spectrum. We briefly address this point in Section 7.2.
@@ 5 Comparison with The Analytical Model
Although the model is designed to describe flows with highly relativistic velocities ($\Gamma\gg 1)$, we perform comparisons across all models, including mildly relativistic shocks with $\Gamma_{u}=2$. The comparison is conducted by fitting the Lorentz factor profile, expressed as a function of the pair-unloaded Thomson optical depth $\tilde{\tau}=2\Gamma_{u}n_{u}\sigma_{T}z$, derived from the analytical model, to the corresponding profile from the simulation through the optimization of the parameters $a$ and $\eta$. [^9] To facilitate the comparison, the optical depth used in the analytical model must be translated into $\tilde{\tau}$. The method for this translation, which is essentially identical to that adopted in GNL18 with a minor modification, is provided in Appendix C. These values depend on the energy and angular distributions of pairs and photons and are expected to be on the order of unity. Accordingly, we conduct the fitting procedure within predefined ranges, with $a$ set to vary between 0.5 and 3, and $\eta$ from 0.25 to 1.0.101010Even if we succeed in obtaining a good fit with values far beyond the predefined range, the solution should be considered unphysical. To determine the best-fit values, we systematically scanned these ranges in increments of $0.05$ for both parameters. The comparison is displayed in Figure 6. It is worth noting that $\tilde{\tau}$ represents the pair-unloaded optical depth measured in the upstream rest frame. Therefore, when a shock propagates in a wind-like environment, the fraction of shock energy released at a given optical depth can be determined by identifying $f_{esc}$ that yields a shock structure with a depth $\tilde{\tau_{u}}(=2\Gamma_{u}\sigma_{T}z_{u})$ comparable to the optical depth.
@@
Figure 6: Comparison of the Lorentz factor profile as a function of pair-unloaded optical depth, $\tilde{\tau}=2\Gamma_{u}n_{u}\sigma_{T}z$, with the analytical model of GNL18 for $\Gamma_{u}=2$ (top left), $3.5$ (top right), 66 (bottom left), and $10$ (bottom right). For the shock with $\Gamma_{u}=2$, the analytical model reproduces the numerical results, with the optimal parameter pairs $(a,\eta)$ for each escape fraction $f_{esc}$ are determined as follows: at $f_{esc}=0\%$, $a=0.55$ and $\eta=0.4$; at $0.55~\%$, $a=0.6$ and $\eta=0.4$; at $f_{esc}=1.5~\%$, $a=0.65$ and $\eta=0.4$; at $f_{esc}=6~\%$ $a=0.7$ and $\eta=0.4$; at $f_{esc}=13~\%$ $a=0.8$ and $\eta=0.4$; at $f_{esc}=35~\%$ $a=1.95$ and $\eta=0.35$. Although the best fit values of the free parameters vary among the different escape fractions, a single pair, e.g., $a=0.5$ and $\eta=0.3$ exhibits a good agreement with the numerical results. On the other hand, for the shock with $\Gamma_{u}=3.5$, the optimal parameter pairs $(a,\eta)$ for each escape fraction $f_{esc}$ are as follows: at $f_{esc}=0\%$, $a=0.5$ and $\eta=0.45$; at $0.45\%$, $a=0.85$ and $\eta=0.4$; at $3.2\%$, $a=0.75$ and $\eta=0.3$; at $7.6\%$, $a=0.75$ and $\eta=0.25$; at $17\%$, $a=0.5$ and $\eta=0.25$; and at $49\%$, $a=0.5$ and $\eta=0.25$. For the high–escape-fraction cases at $\Gamma_{u}=3.5$ ($f_{esc}\geq 0.17$), both parameters sit at the lower bounds of the fitting range ($a=0.5$, $\eta=0.25$), indicating the analytical model’s limitations in capturing the simulation results at high $f_{esc}$. For the shock with $\Gamma_{u}=6$, the optimal pairs are: at $f_{esc}=0\%$, $a=2.4$ and $\eta=0.6$; at $0.65\%$, $a=0.7$ and $\eta=0.75$; at $2.9\%$, $a=0.5$ and $\eta=0.35$; and for $6.8\%$, $13\%$, and $46\%$, both parameters reach the lower bounds ($a=0.5$, $\eta=0.25$). For the shock with $\Gamma_{u}=10$, the optimal pairs are: at $f_{esc}=0\%$, $a=1.75$ and $\eta=0.6$; at $0.41\%$, $a=0.5$ and $\eta=0.8$; at $2.5\%$, $a=0.5$ and $\eta=0.35$; and for $5.9\%$, $14\%$, and $41\%$, both parameters again take the lower bounds ($a=0.5$, $\eta=0.25$). Taken together, the high-$f_{esc}$ points for $\Gamma_{u}=3.5$, $6$, and $10$ saturate at the lower bounds $(a=0.5,\ \eta=0.25)$, consistent with the above indication of the analytical model’s limitations in this regime.
@@ 5.1 Comparison of Shock Profiles
In the case of infinite shocks, we find that, within a reasonable range of $\eta$ and $a$, the analytical model reproduces the numerical results quite well in all cases ($\Gamma_{u}=2$, $3.5$, 66, and $10$). This agreement was previously reported in ILN20a for $\Gamma_{u}=6$ and $10$ (as well as for $\Gamma_{u}=20$, which is not included in the present paper), confirming the model’s validity for highly relativistic shocks. Here, we extend this finding to mildly relativistic shocks ($\Gamma_{u}<6$), showing that even for $\Gamma_{u}=2$, the analytical model exhibits good agreement with the numerical results, demonstrating its applicability to lower-velocity shocks than previously expected. A more comprehensive comparison between the analytical model and the numerical results for infinite shocks is presented in Appendix C. For mildly relativistic shocks with $\Gamma_{u}=2$, we also find a good match in the case of finite shocks, regardless of the value of $f_{esc}$. By contrast, for higher Lorentz factor shocks ($\Gamma_{u}\geq 3.5$), good agreement persists at small $f_{\rm esc}$, whereas a notable discrepancy emerges and grows as $f_{\rm esc}$ increases. This discrepancy cannot be reconciled by varying the parameters $\eta$ and $a$ within physically reasonable ranges; it becomes larger at higher $\Gamma_{u}$. This trend of increasing discrepancy with $\Gamma_{u}$ is also confirmed by our additional two $\Gamma_{u}=15$ runs at $f_{\rm esc}=0.14$ and $0.4$, which are not shown in Figure 6.
@@
The result is counterintuitive, as one would naively expect highly relativistic shocks to match better with the analytical model, given that it is constructed under the assumption of $\Gamma\gg 1$. This discrepancy can be understood as a consequence of the role of the subshock in highly relativistic shocks, which influences the spectral shape of counterstreaming photons. As shown in Section 3, the subshock produces a high-temperature spike in the immediate downstream region, where relativistically heated pairs scatter photons to higher energies. While the post-subshock temperature remains in the range $kT_{d,sub}\lesssim 0.5m_{e}c^{2}$ for $\Gamma_{u}=2$ (at least up to $f_{esc}\leq 35\%$), it exceeds $\sim m_{e}c^{2}$ and continues to rise as the escape fraction increases for $\Gamma_{u}\geq 3.5$. As a result of this high temperature, a high-energy photon component ($h\nu\sim 3kT_{d,sub}\gtrsim 3m_{e}c^{2}$) emerges in addition to the bulk of photons with energy $\sim m_{e}c^{2}$, as described in Section 4. The high-energy extension observed in the escaping photon spectrum above the spectral peak $E_{p}\sim m_{e}c^{2}$ (Figure 5) corresponds to the fraction of these high-energy photons that counterstream up to the upstream boundary. As mentioned in Section 4, these high-energy photons are efficiently absorbed via $\gamma$-$\gamma$ attenuation before reaching the upstream boundary. This absorption of high-energy counterstreaming photons is responsible for the deviation of the numerical results from the analytical model, which can be explained as follows.
@@
The key point is that the absorption of counterstreaming high-energy photons primarily occurs via collisions with the bulk population of lower-energy photons that are likewise counterstreaming upstream from the immediate downstream region, rather than with photons advecting with the flow (i.e., backscattered photons). This occurs because the high-energy photons possess sufficient energy to exceed the pair production threshold necessary for interaction with the bulk of the counterstreaming photon population. As a result, the effective target photon population for the high-energy photons is significantly more abundant than the backscattered photons assumed in the analytical model. As the escape fraction increases, the subshock strengthens, and this channel of $\gamma-\gamma$ pair production becomes the primary mechanism governing the shock deceleration process. Consequently, as $f_{\rm esc}$ increases, the assumption in the analytical model that shock dissipation is governed by the interaction of counterstreaming photons ($h\nu\sim m_{e}c^{2}$) with pairs and backscattered photons begins to break down, and the shock deceleration width becomes significantly shorter due to the high opacity arising from $\gamma$-$\gamma$ interactions among counterstreaming photons. As noted earlier, for $\Gamma_{u}\geq 3.5$ the deviation from the analytical model increases with $\Gamma_{u}$—–a trend we confirm up to $\Gamma_{u}=15$ in the present study. The larger deviations at higher $\Gamma_{u}$ are due to the increasingly pronounced subshock temperature spike, $kT_{d,{\rm sub}}$.
@@ 5.2 Comparison of Shock Widths
In Figure 7, we present the deceleration length of finite shocks, $\Delta\tilde{\tau}$, as a function of $f_{esc}$, compared with predictions from the analytical model. For the free parameters of the analytical model, we adopt the pair of $a$ and $\eta$ that best fit the $\Gamma-\tilde{\tau}$ profile of the numerical results for the lowest $f_{esc}$ cases (profiles shown by green lines in Figure 6). As a reference, we also show two magenta points from simulations with $\Gamma_{u}=15$ in the $\Gamma_{u}=10$ panel. It is worth noting that although the analytical model considered in this study is modified from the original by shifting the reference position of $\tilde{\tau}=0$ to the location where $\Gamma=\Gamma_{u,sub}$—rather than $\Gamma=1$, as in the original analytical solution—this adjustment, while shortening the shock width by removing part of the upstream region, has a negligible impact on its overall value. This is because the shock width is primarily determined by the far upstream region, where $\Gamma\sim\Gamma_{u}$, as evident from Figure 6.
@@
Figure 7: Comparison of the shock width, $\Delta\tilde{\tau}$, obtained from the simulation (blue circles) and the analytical model (black dashed lines) as a function of $f_{esc}$ for $\Gamma_{u}=2$ (top left), $3.5$ (top right), 66 (bottom left), and $\Gamma_{u}=10$ (bottom right). The shock width, $\Delta\tilde{\tau}$, is defined as the optical depth measured from the subshock to the upstream position where the flow’s four-velocity reaches 90% of its far-upstream value ($\Gamma\beta=0.9\Gamma_{u}\beta_{u}$). For the analytical model, the parameter pair $(a,\eta)$ is fixed to the values that best fit the lowest-$f_{esc}$ simulation in each panel. Orange solid lines (shown for $\Gamma_{u}=3.5,6,$ and $10$) indicate the best-fit power law $\Delta\tilde{\tau}=Cf_{\rm esc}^{X}$ to the simulation data in the corresponding panel, obtained by excluding the single point with $f_{esc}<0.05$. Magenta circles in the bottom right panel denote reference points from the $\Gamma_{u}=15$ simulations (shown for comparison only and excluded from the fits).
@@
Since the shock profile is well reproduced by the analytical model, the dependence of the shock width on $f_{esc}$ observed in the simulation is also well captured by the analytical model for $\Gamma_{u}=2$. On the other hand, reflecting the breakdown of the analytical model at high escape fractions, the $\tilde{\tau}-f_{esc}$ dependence observed in the simulation cannot be reproduced by the analytical model for $\Gamma_{u}\geq 3.5$. Note that varying the free parameters does not alleviate this discrepancy, as evidenced by the fact that large values of $f_{esc}$ cannot be described within the framework of the analytical model, as mentioned above. At large escape fractions ($f_{esc}\gg\Gamma_{u}^{-2}$), the analytical model predicts a shock width approximately given by $\Delta\tilde{\tau}\sim\frac{m_{e}}{m_{p}}\frac{\Gamma_{u}}{f_{esc}}=5.4\times 10^{-3}\Gamma_{u,1}f_{esc}^{-1}.$ In contrast, power-law fits of the form $\Delta\tilde{\tau}\propto f_{esc}^{X}$ at fixed $\Gamma_{u}$ yield $X\simeq-1.15$, 1.56-1.56, and 1.93-1.93 for $\Gamma_{u}=3.5$, $6$, and $10$, respectively, when using the three points with $f_{esc}>0.05$; for $\Gamma_{u}=15$, the two available points with $f_{esc}>0.1$ give $X\simeq-2.20$. This indicates a systematically steeper dependence on $f_{esc}$, with the power-law index becoming more negative as $\Gamma_{u}$ increases (i.e., $|X|$ grows). Thus, the simulation exhibits a much narrower shock width at large $f_{esc}$, with the gap between the simulation and the analytical model exceeding an order of magnitude at the highest escape fractions.